OpenStudy (anonymous):

Solve |2 - 3x| = 7. A.{-5/3, 3} B.{-15, 3} C.{-3, 5/3} D.{-3, 15}

OpenStudy (angela210793):

2-3x=7 3x-2=7 so the solution is A -3x=5 3x=9 x=-3/5 x=3